# Second-Order Constrained Dynamic Optimization

### v2

@@ 7.1.2 Results
Fig. 4b and Fig. 4c compare the cost and constraint violation. Again, we note that in Fig. 4c, some algorithms do not appear when their trajectories are feasible. AL-based DDPs (AL, PDAL, and ADMM) show rapid cost reduction with constraint violation. They hit the target first and then gradually satisfy the constraints, where ADMM shows the slowest improvement in constraint satisfaction. The other two DDP methods, IP and barrier DDP, can keep the trajectory feasible during optimization. IP DDP’s slowness comes from the nature of the IP method, where the trajectory is biased to follow the central path as in (10). In this experiment, $\log$ barrier DDP is a good option that balances cost reduction and constraint satisfaction. For SQP, as in the previous case, the multiple-shooting DDP is comparable to other DDP methods or even better in constraint satisfaction, but the shingle-shooting method performs poorly.
(a) Single shooting method.
(b) Multiple shooting method.
Figure 1: Results for inverted pendulum with single and multiple shooting formulation.
@@
(a) Single shooting method.
(b) Multiple shooting method result.
Figure 3: Results for quadpend with single and multiple shooting formulation.
Swimmer: Starting from the initial position, with its nose in the origin and its tail straight, the swimmer tries to hit the target $[5,-1]$ with its nose. The initial control sequence is zeros, which keeps the swimmer in the initial place. The results are shown in Fig. 5 and Fig. 6, where the difference in the algorithms performance becomes more evident than in the previous examples. The $\log$ barrier DDP method gets stuck before hitting the target due to the approximation. IP DDP shows slow progress and cannot hit the target. Although single-shooting AL and PDAL DDP can complete the task in a similar order of constraint violation $(10^{-5})$, PDAL achieves a lower cost.
In this example, multiple-shooting DDPs struggled to solve the problem. Specifically, they can hit the target with a significant equality constraint violation (especially dynamics) early in the optimization process. However, to satisfy the constraints, they start to show conservative motions and finally end up staying at the initial position with the initial state at a high cost. This is because the equality constraints are always satisfied if they do not move. To alleviate the issue, we regularize the state part of $Q_{\tilde{u}\tilde{u}}$ similar to (35) and the technique in (Jallet et al., 2022b), which prevent $x$s from moving too far from the current trajectory. The results presented here are obtained with this regularization strategy. This modification works in AL DDP with approximated Hessian, making the swimmer move forward. Multiple-shooting SQP can achieve the task with the lowest constraint violation among all methods that can complete the task. Overall, single-shooting AL, PDAL DDPs, and multiple-shooting SQP can complete the task with a similar order of constraint violation. PDAL single-shooting DDP achieves the best cost, followed by AL single-shooting DDP, and multiple-shooting SQP. We deduce that the lower cost in DDP is due to the less constrained search space, which contrasts with SQP. Linearized constraints explicitly constrain the search space of SQP, while DDPs capture information implicitly through the $Q$ function.
(a) Initial trajectory.
(b) Cost Reduction
(c) Constraint Violation
(a) Single shooting method result for swimmer.
(b) Multiple shooting method result for swimmer.
Figure 5: Results for swimmer with single and multiple shooting formulation. The initial, middle, and terminal configurations of the swimmer are drawn in red, green, and black.
(a) Cost Reduction
(b) Constraint Violation
7.2 (R2) How robust is each algorithm to varying initial conditions?
To examine robustness while varying the initial conditions, we use the quadpend system and initialize the algorithms with ten different initial hovering trajectories with the pendulum down. The hovering trajectory is achieved by an initial control sequence $u_{1:N-1}=0.5(m_{\rm{q}}+m_{\rm{p}})g_{0}[1,1]^{\mathsf{T}}$, where $m_{\rm{q}}$, $m_{\rm{p}}$, and $g_{0}$ are the mass of the quadrotor part, the mass of the pendulum part, and the gravitational acceleration, respectively. The algorithms solve the same task as in Section 7.1. Tables 3 and 4 present the results of the single- and multiple-shooting algorithms, respectively. A successful run is defined as the quadpend hitting the target. With $m$ and $\sigma$, we denote the mean and standard deviation of the values specified by their subscripts $J$ and $I$, which denote the cost and inequality constraint violation. When all trajectories are feasible, the symbol $\checkmark$ is used. The arrows in the table represent the preferred value. For example, we have $\uparrow$ next to the success rate because a robust algorithm can hit many targets. To provide an overview of the experiment, we show the resulting trajectories in Fig. 7 obtained by multiple-shooting SQP.
@@
Among single-shooting methods, PDAL DDP achieves the lowest mean and standard deviation in cost. Although its constraint violation is not as small as that of AL DDP, it is sufficiently small. The $\log$ barrier and IP DDP methods achieve strict feasibility but have higher costs than AL and PDAL. ADMM has a large constraint violation in the mean cost caused by two infeasible trajectories. Overall, all single-shooting DDP methods are robust to the initial condition. Among the multiple-shooting methods, SQP performs best, achieving the best value on almost all items. We conclude from these results that when constraint violation is the most important factor, $\log$ barrier DDP is the best method. However, if users can accept small constraint violation, single-shooting PDAL DDP or multiple-shooting SQP can achieve lower costs.
(a) Typical failure mode with hovering initial condition.
(b) Initial and Optimal Trajectories with Multiple shooting method.
@@
One key advantage of multiple-shooting methods over single-shooting ones is that they can enjoy good initial guesses, which was not demonstrated in previous experiments. In the multiple-shooting formulation, an arbitrary state sequence can be used as an initial state trajectory as explained in Sections 4.2.3 and 5.2, which can help guide the algorithm away from poor local minima. To showcase this ability, we test multiple-shooting algorithms with a new task where the quadpend flies through a narrow gap of obstacles and reaches a target behind them with ten different initial points. When initialized with hovering states and controls at a single starting point, as in the previous experiment, the performance of all methods is decremented. They cannot complete the task, i.e., hitting the target with a significant constraint violation, and getting stuck before hitting it, etc. We show typical failure trajectories in Fig. 8a.
Table 3: Comparison of single-shooting algorithms initialized with hovering controls at ten different initial states.
@@
Table 4: Comparison of multiple-shooting algorithms initialized with hovering controls at ten different initial states.
@@
$m_{J}$downarrow 29.4 28.3 29.7 28.7
$\sigma_{J}$downarrow 4.37 3.42 3.94 3.28
@@
To circumvent this problem, we could solve a sequence of subproblems with intermediate targets and a short time horizon so that the quadpend can detour rather than get stuck at local minima. However, using the multiple shooting method with an informed initial guess, the problem should be solved without solving subproblems because the trajectory can be biased towards the right solution. To verify this idea, we initialize the multiple shooting algorithms with a state sequence, one of which is shown in the left figure of Fig. 8b. This trajectory is obtained by linear interpolation of four points such that the state trajectory does not hit obstacles. This initial guess leads to a reasonable solution presented in the right figure in Fig. 8b. The same strategy as this interpolation is used for all ten different initial points to obtain informed initial trajectories. The control sequence is initialized with the hover sequence presented in 7.2, which means that the equality constraints are violated on the initial trajectory.
@@ 8 Conclusion
In this paper, we have reviewed two main methods for constrained dynamic optimization: constrained DDP derived based on NLP techniques and SQP for dynamical systems. We have also discussed two distinct representations of these methods, namely, single- and multiple-shooting formulations. In addition, we derived single-shooting PDAL DDP and added it to the comparison. If explicit feedback is not necessary and the state of the system is not substantially large, then multiple-shooting SQP is the most stable method that can achieve the lowest constraint violation. Another advantage of multiple-shooting SQP is that it can accept good initial guesses that help guiding the optimization process to the desired trajectory. On the contrary, DDP methods are preferred when the dimension of the system is large and feedback gains might be required. The single-shooting DDP variant is the most efficient method in terms of computational complexity. When feasibility is more important, the $\log$ barrier DDP variant can be helpful because it can perform optimization while keeping the trajectory always feasible. On the other hand, in cases where cost reduction is more important, AL and PDAL DDP are the best candidate algorithms to use. These two methods also accept an informed initial guess in their multiple-shooting formulations. Finally, we analyzed and experimentally confirmed that PDAL DDP performs better than AL DDP, although the difference was not significant when using GN approximation in multiple-shooting case.
{acks}
This work is supported by NASA under the ULI grant 80NSSC22M0070. Augustinos Saravanos acknowledges support by the A. Onassis Foundation Scholarship.
@@ A.1 General form of SQP
The derivation of this section is based on those of the SQP solver SNOPT Gill et al. (2002) and NPSOL Gill et al. (1986a). Nevertheless, we believe this section is important because some of the important techniques, for instance, solutions to problems to obtain parameters during optimization, were not explicitly explained in these two papers.
@@ A.1.1 Problem formulation of SQP
Consider the constraint optimization problem
$\min_{x}\ f_{0}(x),\quad{\text{s.t.}}\quad c(x)\geq 0,$
where $f_{0}\in\mathbb{R}$ is an objective as in the main article and $c(x)\in\mathbb{R}^{w}$ is a constraint function. We have used $g(x)\leq 0$ for the constraint in the main article, but respecting the existing work for SQP, we use $c(x)=-g(x)\geq 0$ in this section. SQP transforms the original problem into a sequence of Quadratic Programming (QP) with linearized constraints around the current $x$. We use $x_{k}$ to denote the value of $x$ in the current $k$ th iteration. The QP subproblem is given by
${}\min_{\delta x_{k}}\Big{[}f_{0}(x_{k})+[\nabla f_{0}(x_{k})]^{\mathsf{T}}\delta x_{k}+\frac{1}{2}\delta x_{k}^{\mathsf{T}}H\delta x_{k}\Big{]}$ (66)
[…]
where $H$ is Hessian of Lagrangian $\mathcal{L}$
$\mathcal{L}=f_{0}(x)-\lambda^{\mathsf{T}}c(x),\ H=\nabla_{xx}\mathcal{L}$
with nonnegative Lagrangian multiplier $\lambda\in\mathbb{R}^{w}$. Notice that the objective in (66) has a gradient of $f_{0}$ and Hessian of $\mathcal{L}$, not Hessian of $f_{0}$. This discrepancy of gradient and Hessian can be explained using a modified Lagrangian $\mathcal{L}_{\rm{m}}$ (Robinson, 1972; Van Der Hoek, 1982);
$\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})=f_{0}(x)-\lambda_{k}^{\mathsf{T}}d_{\rm{L}}(x,x_{k}),$ (67)
@@
$\nabla d_{\rm{L}}(x,x_{k})=\nabla c(x)-\nabla[c_{\rm{L}}(x,x_{k})]=\nabla c(x)-\nabla c(x_{k}),$
@@
$\nabla\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})$ $=\nabla f_{0}(x)-[\nabla c(x)-\nabla c(x_{k})]^{\mathsf{T}}\lambda_{k},$
$\nabla_{xx}\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})$ $=\nabla_{xx}f_{0}(x)-\sum_{i=1}^{w}\big{[}[\lambda_{k}]_{i}\nabla_{xx}c_{i}(x)\big{]}.$
Evaluated at $x=x_{k}$, $\mathcal{L}_{\rm{m}}$ takes the same value as the objective as
$\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=f_{0}(x_{k}).$
Moreover, its gradient at the same point is equal to that of the objective
$\nabla\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=\nabla f_{0}(x_{k}).$
@@
$\nabla_{xx}\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=\nabla_{xx}f_{0}(x_{k})-\sum_{i=1}^{w}\big{[}[\lambda_{k}]_{i}\nabla_{xx}c_{i}(x_{k})\big{]}.$
Therefore, quadratic approximation of $\mathcal{L}_{\rm{m}}$, at $x=x_{k}$, denoted by $\mathcal{L}_{\rm{mq}}$, is obtained by
@@
$=$ $\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})+[\nabla\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})]^{\mathsf{T}}\delta x_{k}$
$\hskip 42.67912pt+\frac{1}{2}\delta x_{k}^{\mathsf{T}}\nabla_{xx}\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})\delta x_{k}$
@@
with $\delta x_{k}=x-x_{k}$. Given $x_{k}$, minimization of $\mathcal{L}_{\rm{mq}}$, i.e., quadratic approximation of modified Lagrangian, is equivalent to minimization of
@@
which recovers (66). More explanation is found in Boggs and Tolle (1995). After solving the QP subproblem, $x$ is updated by the solution of (66) denoted by $\delta x^{\ast}_{k}$ with step size $\alpha\in(0,1]$ as in (3) where the new $x$ used in the next $k+1$ th iteration. SQP repeats this process iteratively, until KKT condition of (66) given in (26) is satisfied with the given tolerance.

### v3

@@ 7.1.2 Results
Fig. 4(b) and Fig. 4(c) compare the cost and constraint violation. Again, we note that in Fig. 4(c), some algorithms do not appear when their trajectories are feasible. AL-based DDPs (AL, PDAL, and ADMM) show rapid cost reduction with constraint violation. They hit the target first and then gradually satisfy the constraints, where ADMM shows the slowest improvement in constraint satisfaction. The other two DDP methods, IP and barrier DDP, can keep the trajectory feasible during optimization. IP DDP’s slowness comes from the nature of the IP method, where the trajectory is biased to follow the central path as in (10). In this experiment, $\log$ barrier DDP is a good option that balances cost reduction and constraint satisfaction. For SQP, as in the previous case, the multiple-shooting DDP is comparable to other DDP methods or even better in constraint satisfaction, but the shingle-shooting method performs poorly.
Figure 1: Results for inverted pendulum.
Figure 3: Results for quadpend.
Swimmer: Starting from the initial position, with its nose in the origin and its tail straight, the swimmer tries to hit the target $[5,-1]$ with its nose. The initial control sequence is zeros, which keeps the swimmer in the initial place. The results are shown in Fig. 5 and Fig. 6, where the difference in the algorithm’s performance becomes more evident than in the previous examples. The $\log$ barrier DDP method gets stuck before hitting the target due to the approximation. IP DDP shows slow progress and cannot hit the target. Although single-shooting AL and PDAL DDP can complete the task in a similar order of constraint violation $(10^{-5})$, PDAL achieves a lower cost.
In this example, multiple-shooting DDPs struggled to solve the problem. Specifically, they can hit the target with a significant equality constraint violation (especially dynamics) early in the optimization process. However, to satisfy the constraints, they start to show conservative motions and finally end up staying at the initial position with the initial state at a high cost. This is because the equality constraints are always satisfied if they do not move. To alleviate the issue, we regularize the state part of $Q_{\tilde{u}\tilde{u}}$ similar to (35) and the technique in (Jallet et al., 2022b), which prevent $x$s from moving too far from the current trajectory. The results presented here are obtained with this regularization strategy. This modification works in AL DDP with approximated Hessian, making the swimmer move forward. Multiple-shooting SQP can achieve the task with the lowest constraint violation among all methods that can complete the task. Overall, single-shooting AL, PDAL DDPs, and multiple-shooting SQP can complete the task with a similar order of constraint violation. PDAL single-shooting DDP achieves the best cost, followed by AL single-shooting DDP, and multiple-shooting SQP.
Panda: The task of the arm is to place the end effector on the target position and stop, while avoiding four spherical obstacles, and satisfying the joint and its command limit. In this example, the second-order information of dynamics is not stably available from the simulator. Therefore, we drop the corresponding terms of constraints in AL and PDAL multiple-shooting DDPs. Consequently, the AL multiple-shooting method only has an approximate version. The initial position of the arm is given by the joint angles from the base link as $x_{0}=[0,0,0,-\pi/2,0,\pi/2,-\pi/4]$. The initial control command maintains the initial state. The results are presented in Fig. 7 and Fig 8. The algorithms find different local solutions, as in the quadpend. Similarly to the swimmer example, $\log$-barrier and IP DDP get stuck and cannot hit the target, while other single-shooting DDPs can complete the task. An addition of the feasibility restoration mechanism could improve the performance of IP DDP. AL and PDAL show similar performance, while ADMM has high constraint violation. This is because we use a small penalty parameter to prioritize hitting the target rather than satisfying constraints in ADMM DDP. For multiple-shooting DDPs, although both AL and PDAL can reduce constraint violations over iterations, only PDAL can achieve the task. SQPs show a tendency similar to that in other experiments. The multiple-shooting one can complete the task with a similar magnitude of constraint violation as other successful DDPs, while the single-shooting one performs poorly.
Figure 5: Results for swimmer. The initial, middle, and terminal configurations of the swimmer are drawn in red, green, and black.
Figure 7: Results for panda. The target and obstacles are shown in red and gray spheres, respectively. The trajectory of the end-effector is drawn in blue spheres.
@@
Figure 8: Comparison of algorithms with a single initial trajectory. Panda.
7.2 (R2) How robust is each algorithm to varying initial conditions and targets?
Initial conditions: To examine robustness while varying the initial conditions, we use the quadpend system and initialize the algorithms with ten different initial hovering trajectories with the pendulum down. The hovering trajectory is achieved by an initial control sequence $u_{1:N-1}=0.5(m_{\rm{q}}+m_{\rm{p}})g_{0}[1,1]^{\mathsf{T}}$, where $m_{\rm{q}}$, $m_{\rm{p}}$, and $g_{0}$ are the mass of the quadrotor part, the mass of the pendulum part, and the gravitational acceleration, respectively. The algorithms solve the same task as in Section 7.1. Tables 3 and 4 present the results of the single- and multiple-shooting algorithms, respectively. A successful run is defined as the quadpend hitting the target. With $m$ and $\sigma$, we denote the mean and standard deviation of the values specified by their subscripts $J$, $I$, and $E$, which denote the cost, inequality, and equality constraint violation. When all trajectories are feasible, the symbol $\checkmark$ is used. The arrows in the table represent the preferred value. For example, we have $\uparrow$ next to the success rate because a robust algorithm can hit many targets. To provide an overview of the experiment, we show the resulting trajectories in Fig. 9 obtained by multiple-shooting SQP.
@@
Among single-shooting methods, PDAL DDP achieves the lowest mean and standard deviation in cost. Although its constraint violation is not as small as that of AL DDP, it is sufficiently small. The $\log$ barrier and IP DDP methods achieve strict feasibility but have higher costs than AL and PDAL. ADMM has a large constraint violation in the mean cost caused by two infeasible trajectories. Overall, all single-shooting DDP methods are robust to the initial condition. Among the multiple-shooting methods, SQP performs best, achieving the best value on almost all items.
@@
Figure 11: Results of Panda for multiple targets.
(a) Single-shooting PDAL DDP
(b) Multiple-shooting SQP
Figure 12: Comparison of a failed trajectory of single-shooting PDAL DDP and a successful trajectory from multiple-shooting SQP.
Different targets: To evaluate the robustness of the algorithms across different targets and tasks, we design three obstacle fields, containing one, two, and four obstacles, respectively. For each field, we conduct one, four, and five experiments with varying target configurations. In total, we have ten experiments. We note that one of these environments, i.e, a pair of an obstacle field and a target, is the same as the one used in the previous example. The results are provided in Table 5 and Table 6. The overview of the experiments is shown in Fig. 11.
In single-shooting methods, $\log$ barrier and IP DDPs show conservative behavior to take distance from constraints through the barrier function. As a result, their performance decreases in cluttered environments. Among the single-shooting algorithms, PDAL DDP has the highest success rate with sufficiently small constraint violation. The multiple-shooting SQP outperforms it in terms of success rate with equality constraint violation. In Fig. 12, we compare a failed trajectory from single-shooting PDAL DDP and a successful trajectory from multiple-shooting SQP in the same task. In the early time steps, both of them draw similar trajectories. However, only SQP can successfully hit the target by rotating a joint close to the base. In contrast, a local minimum captures PDAL DDP, where it continues to bend the arm without incorporating the rotation observed in SQP. Panda can locate the end effector close to the target in this local solution, but cannot hit it. Although single-shooting DDP failed in this example, it can achieve a lower mean cost compared to multiple-shooting SQP.
We conclude from these results that when constraint violation is the most critical factor, $\log$ barrier DDP is the best method. However, it might not be able to complete the task, especially in high-dimensional systems and cluttered environments. If users can accept small constraint violations, single-shooting PDAL DDP or multiple-shooting SQP has a higher success rate. In our example, SQP shows a slightly better success rate, whereas PDAL DDP achieves a lower mean cost. A key difference between these two algorithms is that PDAL DDP can always satisfy dynamics, whereas SQP has small violations.
@@
One key advantage of multiple-shooting methods over single-shooting ones is that they can enjoy good initial guesses, which was not demonstrated in previous experiments. In the multiple-shooting formulation, an arbitrary state sequence can be used as an initial state trajectory as explained in Sections 4.2.3 and 5.2, which can help guide the algorithm away from poor local minima. To showcase this ability, we test multiple-shooting algorithms with a new task where the quadpend flies through a narrow gap of obstacles and reaches a target behind them with ten different initial points. When initialized with hovering states and controls at a single starting point, as in the previous experiment, the performance of all methods is decremented. They cannot complete the task, i.e., hitting the target with a significant constraint violation, and getting stuck before hitting it, etc. We show typical failure trajectories in Fig. LABEL:fig:quad_fail_through.
Table 3: Comparison of single-shooting algorithms initialized with hovering controls at ten different initial states for quadpend.
@@
Table 4: Comparison of multiple-shooting algorithms initialized with hovering controls at ten different initial states for quadpend.
@@
$m_{J}$downarrow 29.4 28.3 29.7 28.5
$\sigma_{J}$downarrow 4.37 3.42 3.94 3.39
[…]
Table 5: Comparison of single-shooting algorithms for ten different targets for Panda.
Log Barrier AL single IP PDAL single ADMM single SQP single
Success rate [%] $\uparrow$ 60 80 50 90 90 60
$m_{J}$downarrow 7.94 8.81 8.80 8.82 9.41 10.7
$\sigma_{J}$downarrow 1.50 2.54 1.99 2.48 2.67 3.28
[…]
Table 6: Comparison of multiple-shooting algorithms for ten different targets for Panda.
AL multi PDAL multi SQP multi
Success rate [%] $\uparrow$ 0 60 100
$m_{J}$downarrow - 10.3 9.09
$\sigma_{J}$downarrow - 2.36 2.45
[…]
@@
To circumvent this problem, we could solve a sequence of subproblems with intermediate targets and a short time horizon so that the quadpend can detour rather than get stuck at local minima. However, using the multiple shooting method with an informed initial guess, the problem should be solved without solving subproblems because the trajectory can be biased towards the right solution. To verify this idea, we initialize the multiple shooting algorithms with a state sequence, one of which is shown in the left figure of Fig. 10(b). This trajectory is obtained by linear interpolation of four points such that the state trajectory does not hit obstacles. This initial guess leads to a reasonable solution presented in the right figure in Fig. 10(b). The same strategy as this interpolation is used for all ten different initial points to obtain informed initial trajectories. The control sequence is initialized with the hover sequence presented in 7.2, which means that the equality constraints are violated on the initial trajectory.
@@ 8 Conclusion
In this paper, we have reviewed two families of algorithms for constrained dynamic optimization: constrained DDP derived based on NLP techniques and SQP for dynamical systems. We have also discussed two distinct representations of these methods, namely, single- and multiple-shooting formulations. In addition, we derived a novel single-shooting PDAL DDP and added it to the comparison. Working towards our goal to systematize the research on second-order constrained dynamic optimization, we performed extensive benchmarking and analyzed algorithms based on criteria such as objective function minimization, task completion rate, and constraint satisfaction. Among the different methods, the single-shooting PDAL DDP and multiple-shooting SQP algorithms stand out due to their consistent performance and robust numerical behavior across different systems and tasks. Both algorithms handle the high-dimensionality and non-convexity of trajectory optimization tasks in robotics well. When the state of the system is not large and a small violation of dynamics is allowed, multiple-shooting SQP is the most stable method that can achieve the highest success rate of tasks. Another advantage of the method is that it can accept good initial guesses that help guide the optimization process to the desired trajectory. On the other hand, when the state of the system is large, and the infeasibility of dynamics is not allowed, the single-shooting PDAL DDP is suitable. This is because the method inherits the low computational complexity and dynamical feasibility of the single-shooting DDP. The method has similar performance to the single-AL variant when the problem is simple, but starts to show its superiority as the problem becomes more complex. These two AL-based DDP methods can also be formulated in multiple-shooting formulations, which can take advantage of informed initial guesses. Finally, we would like to note that although $\log$ barrier DDP struggles to complete tasks, it can be the algorithm of choice when feasibility is prioritized.
Author Contributions
Yuichiro Aoyama: Conceptualization, Formal analysis, Investigation, Software, Methodology, Writing-original draft. Oswin So: Conceptualization, Formal analysis, Investigation, Methodology, Software, Writing-review & editing. Augustinos Saravanos: Investigation, Methodology, Writing-review & editing. Evangelos Theodorou: Conceptualization, Supervision, Project administration, Writing-review & editing.
Statements and Declarations
Ethical considerations
This article does not contain any studies with human or animal participants.
Consent to participate
This article does not contain any studies with human or animal participants.
Consent for publication
Not applicable. {dci} The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article. {funding} The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: Evangelos Theodorou was supported by NASA [grant number 80NSSC22M0070]; Augustinos Saravanos was supported by the A. Onassis Foundation Scholarship.
Appendix
@@ A.1 General form of SQP
The derivation of this section is based on those of the SQP solver SNOPT (Gill et al., 2002) and NPSOL (Gill et al., 1986a). Nevertheless, we believe this section is important because some of the important techniques, e.g., solutions for problems to obtain parameters during optimization, were not explicitly explained.
@@ A.1.1 Problem formulation of SQP
Let us revisit a SQP subproblem in (25). Notice that the objective in has a gradient of $f_{0}$ and Hessian of $\mathcal{L}$, not Hessian of $f_{0}$. This discrepancy can be explained using a modified Lagrangian $\mathcal{L}_{\rm{m}}$ (Van Der Hoek, 1982; Robinson, 1972). Here we write $x=x_{k}+\delta x_{k}$ with a subscript for iteration.
$\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})=f_{0}(x)+\lambda_{k}^{\mathsf{T}}d_{\rm{L}}(x,x_{k}),$ (66)
@@
$\nabla d_{\rm{L}}(x,x_{k})=\nabla g(x)-\nabla[g_{\rm{L}}(x,x_{k})]=\nabla g(x)-\nabla g(x_{k}),$
@@
$\nabla\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})$ $=\nabla f_{0}(x)+[\nabla g(x)-\nabla g(x_{k})]^{\mathsf{T}}\lambda_{k},$
$\nabla_{xx}\mathcal{L}_{\rm{m}}(x,x_{k},\lambda_{k})$ $=\nabla_{xx}f_{0}(x)+\sum_{i=1}^{w}\big{[}[\lambda_{k}]_{i}\nabla_{xx}g_{i}(x)\big{]}.$
Evaluated at $x=x_{k}$, $\mathcal{L}_{\rm{m}}$ and its gradient are equal to those of thr objective as
$\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=f_{0}(x_{k}),\ \nabla\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=\nabla f_{0}(x_{k}).$
@@
$\nabla_{xx}\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k})=\nabla_{xx}f_{0}(x_{k})+\sum_{i=1}^{w}\big{[}[\lambda_{k}]_{i}\nabla_{xx}g_{i}(x_{k})\big{]}.$
Therefore, quadratic approximation of $\mathcal{L}_{\rm{m}}$, at $x=x_{k}$, denoted by $\mathcal{L}_{\rm{mq}}$, is obtained as
@@
Given $x_{k}$, minimization of $\mathcal{L}_{\rm{mq}}$ is equivalent to that of
@@
which recovers (25). More explanation is found in Boggs and Tolle (1995). After solving the QP subproblem, $x_{k}$ is updated by the solution of (25) denoted by $\delta x^{\ast}_{k}$, giving $x_{k+1}$ for a new subproblem.